A square coil RSTU of sides 5.0 cm carrying a current, I of 2.5 A is placed in a uniform magnetic field, B of 2.4 T.Calculate the torque on the coil.


1
Expert's answer
2022-03-14T13:06:58-0400

We know that

Magnetic moment

"M=NIA"

"M=NIa^2"


"M=1\times2.5\times0.05^2=6.25\times10^{-3}Am^2"

Torque

"\tau=M\times B"

"\tau=MBsin\theta"

"\tau=MBsin90°"


"\tau=MB=6.25\times10^{-3}\times2.4=0.015Nm"